The lengths of the sides of a triangle are consecutive odd intergers. If the perimeter is 1 less than 4 times the shortest side, find the length of each side.
Thank you!

Answer by mananth(16949)

The lengths of the sides of a triangle are consecutive odd intergers. If the perimeter is 1 less than 4 times the shortest side, find the length of each side.

let the length of the sides be x, x+2,x+4
..
x+x+2+x+4=4x-1
3x+6=4x-1
4x-3x=6+1
x=7
7, 9, 11
m.ananth@hotmail.ca